OpenStudy (anonymous):

How many grams of ethane gas (C2H6) are in a 12.7 liter sample at 1.6 atmospheres and 24C? Show all work used to solve this problem.

thomaster (thomaster):

\(\Huge\bf \color{black}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{black}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{#ff7100}{Welcome~to~\color{#00B4ff}{Open}}\color{#7cc517}{Study!!}\) Do you know the ideal gas law? \(\large PV=nRT\)

OpenStudy (anonymous):

Not really.

thomaster (thomaster):

P= pressure in atm V= volume inliter n= moles of gas R= gas constant (0.0820574614L  atm K−1 mol−1 T= temperature in kelvin

thomaster (thomaster):

You know the pressure, the volume, the gas constant and the temperature. You only need to calculate n To do that you need to convert the fomula to \(\Large n=\dfrac{PV}{RT}\) 24°C = 297,15 Kelvin \(\Large n=\dfrac{1.6*12.7}{0.0820574614*297.15}\) Then you have the moles of ethane gas. Use the attached diagram to convert to gram.